# Probabilistic Analysis of Copyright Disputes and Generative AI Safety

### v4

@@ Appendix C Proofs of Propositions
Using the logical relationship between copying and access (Definition 3.1 (1)) as well as $EA$’s independence (Definition 3.1 (4)), for any pair $(i,j)$,
$\begin{split}&P\left(C=1|S=i,\,EA=j\right)\\ &=P\left(C=1|A=1,\,S=i\right)P\left(A=1|S=i,\,EA=j\right)\\ &\quad+P\left(C=1|A=0,\,S=i\right)P\left(A=0|S=i,\,EA=j\right)\\ &=P\left(C=1|A=1,\,S=i\right)P\left(A=1|S=i,\,EA=j\right)\\ &\quad+0\cdot P\left(A=0|S=i,\,EA=j\right)\\ &=\alpha_{i}\beta_{i,j}.\end{split}$
@@
The first element of the inverse ratio rule is proved by showing contradiction. Assume $\mu>\mu^{\prime}$ where:
$\begin{gathered}\mu:=\min\{i:P(C=1|S=i,\,EA=j)>\lambda\}\\ \mu^{\prime}:=\min\{i:P(C=1|S=i,\,EA=j^{\prime})>\lambda\}.\end{gathered}$
Then, by assumptions about the ordering of the terms of the form $\alpha_{i}$ and $\beta_{i,j}$,
$\begin{split}\lambda<\alpha_{\mu^{\prime}}\beta_{\mu^{\prime},j^{\prime}}\leq\alpha_{\mu^{\prime}}\beta_{\mu^{\prime},j}\leq\alpha_{\mu}\beta_{\mu,j}\end{split}.$
@@
However, this implies that $\mu^{\prime}$, which is strictly smaller than $\mu$, would have achieved $P(C=1|S=\mu^{\prime},\,EA=j)>\lambda$, contradicting the assumption that $\mu$ was the minimum that achieves this.
@@
The proof for the second element of the inverse ratio rule is similar. Assume $\kappa>\kappa^{\prime}$ where:
$\begin{gathered}\kappa:=\min\{j:P(C=1|S=i,\,EA=j)>\lambda\}\\ \kappa^{\prime}:=\min\{j:P(C=1|S=i^{\prime},\,EA=j)>\lambda\}.\end{gathered}$
Same as above, the following inequalities hold:
$\lambda<\alpha_{i^{\prime}}\beta_{i^{\prime},\kappa^{\prime}}\leq\alpha_{i}\beta_{i,\kappa^{\prime}}\leq\alpha_{i}\beta_{i,\kappa}$
which results in contradiction with the assumption of the minimality of $\kappa$.
C.2. Proof of Proposition 4.2
By the NAF condition (Definition 4.1 (1)) and conditional independence (Definition 4.1 (3)):
@@
By rearranging terms,
$\begin{split}P(A_{M}=1|S=i,\,EA_{M}=k)&\leq\frac{e^{\epsilon}P(A_{M}=1|EA_{M}=k)}{1+P(A_{M}=1|EA_{M}=k)(e^{\epsilon}-1)}\\ &=\frac{e^{\epsilon}\delta_{k}}{1+\delta_{k}(e^{\epsilon}-1)}\\ &=\Gamma(\epsilon,\delta_{k}).\end{split}$
Taking the derivative of $\Gamma(\epsilon,\delta)$ with respect to $\epsilon$ and $\delta$, respectively,
$\begin{split}&\frac{d\Gamma}{d\mathcal{\epsilon}}=\frac{e^{\epsilon}\delta(1-\delta)}{(1+\delta(e^{\epsilon}-1))^{2}}\geq 0\\ &\frac{d\Gamma}{d\delta}=\frac{e^{\epsilon}}{(1+\delta(e^{\epsilon}-1))^{2}}\geq 0.\end{split}$
Therefore, $\Gamma(\epsilon,\delta)$ is increasing in $\epsilon$ and $\delta$.
The inequality $\delta\leq\Gamma(\epsilon,\delta)$ follows from the observation that:
$\Gamma(\epsilon,\delta)-\delta=\frac{\delta(e^{\epsilon}-1)(1-\delta)}{1+\delta(e^{\epsilon}-1)}\geq 0$
where the inequality follows from the fact that each term in the numerator and denominator is non-negative because $1\geq\delta\geq 0$ and $e^{\epsilon}>1$.
Finally, the limit as $\epsilon$ approaches zero can be computed straightforwardly as:
$\lim_{\epsilon\rightarrow 0}\Gamma(\epsilon,\delta)=\frac{1\cdot\delta}{1+\delta\cdot 0}=\delta.$
C.3. Proof of Proposition A.1
As in the proof of Proposition 3.2,
$P(C=1|S=i,\,EA=j)=\alpha_{i}\beta_{i,j},\,\forall(i,j).$
For the first part, observe:
$\lambda<\alpha_{\mu^{\prime}}\beta_{\mu^{\prime},j^{\prime}}\leq\alpha_{\mu^{\prime}-1}\beta_{\mu^{\prime}-1,j}.$
In particular, we have $\lambda<\alpha_{\mu^{\prime}-1}\beta_{\mu^{\prime}-1,j}$. This implies $\min\{i:P(C=1|S=i,\,EA=j)>\lambda\}\leq\mu^{\prime}-1<\mu^{\prime}$, which completes the proof of the first part. Similarly, for the second part, we have:
$\lambda<\alpha_{i^{\prime}}\beta_{i^{\prime},\kappa^{\prime}}\leq\alpha_{i}\beta_{i,\kappa^{\prime}-1}.$
This implies $\min\{j:P(C=1|S=i,\,EA=j)>\lambda\}\leq\kappa^{\prime}-1<\kappa^{\prime}$, which completes the proof.

### v5

@@ Appendix B Proofs
From the logical relationship between copying and access as well as $EA$’s conditional independence, we have $P\left(C=1|S=i,\,EA=j\right)=\alpha_{i}\beta_{i,j}$ for any $(i,j)$.
@@
The first element of the inverse ratio rule is proved by showing contradiction. Assume $\mu>\mu^{\prime}$ where $\mu:=\min\{i:P(C=1|S=i,\,EA=j)>\lambda\}$ and $\mu^{\prime}:=\min\{i:P(C=1|S=i,\,EA=j^{\prime})>\lambda\}$.
@@
Then, by assumptions about the ordering of the terms of the form $\alpha_{i}$ and $\beta_{i,j}$, the inequalities $\lambda<\alpha_{\mu^{\prime}}\beta_{\mu^{\prime},j^{\prime}}\leq\alpha_{\mu^{\prime}}\beta_{\mu^{\prime},j}$ hold. However, this implies that $\mu^{\prime}$, which is strictly smaller than $\mu$, would have achieved $P(C=1|S=\mu^{\prime},\,EA=j)>\lambda$, contradicting the assumption that $\mu$ was the minimum that achieves this.
@@
The proof for the second element of the inverse ratio rule is similar. Assume $\kappa>\kappa^{\prime}$ where $\kappa:=\min\{j:P(C=1|S=i,\,EA=j)>\lambda\}$ and $\kappa^{\prime}:=\min\{j:P(C=1|S=i^{\prime},\,EA=j)>\lambda\}$. Then, the inequalities $\lambda<\alpha_{i^{\prime}}\beta_{i^{\prime},\kappa^{\prime}}\leq\alpha_{i}\beta_{i,\kappa^{\prime}}$ hold, which is in contradiction with the assumption of minimality of $\kappa$. ∎
Proof of Proposition 3.2.
By the NAF condition and the conditional independence of $EA_{M}$:
@@
Rearranging the terms, we obtain the upper bound $\Gamma(\epsilon,\delta_{k})$.
Taking the derivatives of $\Gamma(\epsilon,\delta)$, we get $\frac{\partial\Gamma}{\partial\mathcal{\epsilon}}=\frac{e^{\epsilon}\delta(1-\delta)}{(1+\delta(e^{\epsilon}-1))^{2}}\geq 0$ and $\frac{\partial\Gamma}{\partial\delta}=\frac{e^{\epsilon}}{(1+\delta(e^{\epsilon}-1))^{2}}\geq 0$. Thus, $\Gamma(\epsilon,\delta)$ is increasing in $\epsilon$ and $\delta$. The inequality $\delta\leq\Gamma(\epsilon,\delta)$ follows from the observation that $\Gamma(\epsilon,\delta)-\delta=\frac{\delta(e^{\epsilon}-1)(1-\delta)}{1+\delta(e^{\epsilon}-1)}\geq 0$, which holds since $1\geq\delta\geq 0$ and $e^{\epsilon}>1$. Finally, the limit as $\epsilon$ approaches zero can be computed straightforwardly as: $\lim\limits_{\epsilon\rightarrow 0}\Gamma(\epsilon,\delta)=\frac{1\cdot\delta}{1+\delta\cdot 0}=\delta$. ∎
Proof of Proposition A.1.
As in the proof of Proposition 2.2, $P(C=1|S=i,\,EA=j)=\alpha_{i}\beta_{i,j}$ for any pair $(i,j)$. For the first part, observe: $\lambda<\alpha_{\mu^{\prime}}\beta_{\mu^{\prime},j^{\prime}}\leq\alpha_{\mu^{\prime}-1}\beta_{\mu^{\prime}-1,j}$. In particular, we have $\lambda<\alpha_{\mu^{\prime}-1}\beta_{\mu^{\prime}-1,j}$. This implies $\min\{i:P(C=1|S=i,\,EA=j)>\lambda\}\leq\mu^{\prime}-1<\mu^{\prime}$, which completes the proof of the first part. Similarly, for the second part, we have: $\lambda<\alpha_{i^{\prime}}\beta_{i^{\prime},\kappa^{\prime}}\leq\alpha_{i}\beta_{i,\kappa^{\prime}-1}$. This implies $\min\{j:P(C=1|S=i,\,EA=j)>\lambda\}\leq\kappa^{\prime}-1<\kappa^{\prime}$, which completes the proof. ∎